Temporary Housing Services Incorporated (THSI) is considering a project that involves setting up a temporary housing facility in an area recently damaged by a hurricane. THSI will lease space in this facility to various agencies and groups providing relief services to the area. THSI estimates that this project will initially cost $5.51 million to setup and will generate $20 million in revenues during its first and only year in operation (paid in one year). Operating expenses are expected to total $12 million during this year and depreciation expense will be another $3 million. THSI will require no working capital for this investment. THSI's marginal tax rate is 35%.

Assume that THSI's cost of capital is 16.3% p.a.

Compute the NPV of the temporary housing facility to the nearest dollar. (Do not enter a dollar sign, just enter your answer as a whole number, either positive or negative)
Answer #1

Expcted Revenue = $ 20 million

Less: Operating Expenses = $ 12 million

Gross profit = $ 8 million

Less:Depreciation Expense = $ 3 million

EBIT = $ 5 million

Less: Tax @ 35 % = 0.35 x 5 = $ 1,75 million

NOPAT (Net Operating Profit After-Tax) = $ 3.25 million

FCFF = NOPAT + Depreciation - Capital Expenditure - Investment in Working Capital = 3.25 + 3 - 0 = $ 6.25 million

Initial Investment = $ 5.51 million and Cost of Capital = 16.3 %

Therefore, Project NPV = (6.25 / 1.163) - 5.51 = - 0.135967 million or - 135967